Tomatoes that regularly sell for $0.64 per pound are on sale for 25% off the regular price. What is the sale price?
Answer by ikleyn(53942)

.

0.64 - 0.25*064 = 0.64 - 0.16 = 0.48 dollars per pound.

Solved.

---------------

On percentage,  see the lessons
    - Percentage problems
    - Percentage word problems (Type 1 problems, Finding the Part)
    - Percentage word problems (Type 2 problems, Finding the Rate)
    - Percentage word problems (Type 3 problems, Finding the Base)
    - Simple percentage problems
    - More complicated percentage problems
    - Advanced problems on percentage
    - Percentage problems on chains of discounts
    - Problems on percentage that lead to unexpected results
    - OVERVIEW of lessons on percentage problems
in this site.

They are written specially for you.

Read them and become an expert in this area.

Start from the very first lesson of the list.

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.